# Sparse Activations as Conformal Predictors

### v1

@@ 1 INTRODUCTION
The use of learned predictive models in many high-stakes applications (e.g., medical, legal, or financial) has stimulated extensive research on uncertainty quantification as a way to enhance predictions with reliable confidence estimates (Silva-Filho et al.,, 2023; Gawlikowski et al.,, 2023). In classification tasks, this corresponds to providing an estimate of the posterior probabilities of each of the classes, given the sample to be classified. In regression, this would correspond to providing, not only a point estimate, but also a confidence interval around that estimate.
@@
Unfortunately, modern deep networks are not well calibrated: the class probability estimates may be significantly different from the posterior probability, and thus cannot be trusted as confidence levels (Guo et al.,, 2017; Wenger et al.,, 2020). This observation has motivated considerable work on developing methods to obtain well-calibrated class probability estimates.
@@
To address this problem, conformal prediction2.1) has emerged as a powerful framework for uncertainty quantification that comes with strong theoretical guarantees under minimal assumptions, as it is model-agnostic and distribution-free (Vovk et al.,, 2005; Angelopoulos and Bates,, 2023). Conformal predictors are set predictors, i.e., their outputs are either sets of labels (for classification tasks) or intervals (for regression tasks). Most common variants, namely split conformal prediction (Papadopoulos et al.,, 2002), do not require model retraining: prediction rules are created using a separate calibration set through the design of a non-conformity score, which is a measure of how unlikely an input-output pair is.
@@
A parallel research direction concerns sparse alternatives to the softmax activation (§2.2), i.e., capable of producing probability outputs where some labels receive zero probability, yet are differentiable, thus usable in backpropagation (Martins and Astudillo,, 2016; Peters et al.,, 2019). The class of Fenchel-Young losses (Blondel et al.,, 2020) provides a framework to learn such sparse predictors, generalizing the softmax activation and the cross-entropy loss. An added benefit of sparse probability outputs is that the labels with nonzero probability can be directly interpreted as a set prediction, eliminating the need for creating these sets by thresholding probability values. Such set predictions have been used by Martins and Astudillo, (2016) for multi-label classification, but not in the scope of uncertainty quantification.
@@
This paper connects the two lines of work above by conformalizing sparse activations (Figure 1; §3). We focus on the $\gamma$-entmax family [^1] Called $\alpha$-entmax by Peters et al., (2019). We choose $\gamma$ in this paper to avoid clash with the confidence level $\alpha$, commonly used in the conformal prediction literature. (Peters et al.,, 2019), which includes softmax ($\gamma=1$) and sparsemax ($\gamma=2$) as particular cases, and is sparse for $\gamma>1$. We create set predictions by including all labels receiving nonzero probability under those activations—these sets can be made larger or smaller by controlling a temperature parameter. We leverage the calibration techniques of conformal prediction to choose the temperature that leads to the desired coverage level.
@@ 2.1 Conformal Prediction
We focus on split conformal prediction, which does not involve model retraining (Papadopoulos et al.,, 2002). Given the predictor, $f$, and a collection of calibration samples, $\mathcal{D}_{\text{cal}}=\bigl{(}(x_{1},y_{1}),...,(x_{n},y_{n})\bigr{)}$, a conformal predictor is defined by a non-conformity score $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. The non-conformity score measures how unlikely an input-output pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is, compared to the remaining calibration data. Ideally, predictions $\hat{y}\in\mathcal{Y}$ yielding pairs $(x_{\text{test}},\hat{y})$ that are likely to occur in the data should have a low non-conformity score, and should thus be included in the prediction set $\mathcal{C}_{\alpha}(x_{\mathrm{test}})$.
@@
as proved by Vovk et al., (2005), without any other assumptions on the predictive model or the data distribution. Ideally, we would like the predictor with the required coverage to be efficient, i.e., to have a small average prediction set size $\mathbb{E}[|C_{\alpha}(X)|]$, and adaptive, i.e., instances that are harder to predict should yield larger prediction sets, representing higher uncertainty.
@@ 2.2 Sparse Activations
Since the exponential function is stricly positive, the softmax transformation never outputs any zero. Martins and Astudillo, (2016) introduced an alternative transformation, sparsemax, which can produce sparse outputs while being almost-everywhere differentiable:
@@
As shown by Peters et al., (2019) and Blondel et al., (2020), sparsemax and softmax are particular cases of a family of $\gamma$-entmax transformations, defined as
@@
where $H_{\gamma}$ is a generalized entropy function (Tsallis,, 1988), defined, for $\gamma>0$, as
@@
Peters et al., (2019) show that the solution of (5) has the form
@@
where $\mathbf{1}$ denotes a vector of ones, $\left[x\right]_{+}=\max\{x,0\}$, and $\tau$ is the (unique) constant ensuring normalization. A bisection algorithm for computing $\gamma$-entmax for general $\gamma$, as well as a more efficient algorithm for $\gamma=1.5$, have been proposed by Peters et al., (2019). We make use of these algorithms in our experiments in §4.
@@
Temperature scaling (Guo et al.,, 2017; Platt et al.,, 1999) is a simple and popular technique to calibrate the output probabilities from a predictive model. By multiplying the label scores $\bm{z}$ by a scaling factor $\beta>0$, which can be interpreted as an inverse temperature parameter, the probabilities resulting from the subsequent softmax transformation can be made more or less peaked. When $\beta=0$, $\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ becomes a uniform distribution, whereas in the limit $\beta\rightarrow+\infty$ (the “zero temperature limit”), $\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ tends to a one-hot vector, assigning probability 1 to the label with the largest score. [^4] If there are multiple labels tied with the same largest score, $\lim_{\beta\rightarrow+\infty}\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ becomes a uniform distribution over those labels, assigning zero probability to all other labels.
Analogous limit cases apply also to $\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$, but there is a key qualitative difference (Martins and Astudillo,, 2016; Blondel et al.,, 2020). For any $\gamma>1$, there is a finite threshold $\beta^{*}$ (corresponding to a “non-zero temperature”) above which $\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$ saturates as the zero temperature limit: $\beta\geq\beta^{*}\,\Rightarrow\,\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})=\lim_{\beta\rightarrow+\infty}\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$. Increasing $\beta$, which may be seen as a regularization coefficient, from 00 to $\beta^{*}$ draws a regularization path where the support of $\gamma$-$\operatorname*{\mathsf{entmax}}(\beta\bm{z})$ varies from the full set of labels to a single label (assuming no ties), as shown in Figure 3. Therefore, using temperature scaling with $\gamma$-entmax, for $\gamma>1$, allows the sparsity of the output to be controlled. We exploit this fact in our proposed method described in §3.
@@ 3.2 Conformalizing $\gamma$-entmax
which is the log-odds ratio between the most probable class and the true one. Since the $\log$ function is monotonic, calibration of this non-conformity score leads to thresholding the odds ratio $p_{1}/p_{k(y)}$, which has an intuitive interpretation: labels whose probability is above a fraction of that of the most probable label are included in the prediction set. [^6] This non-conformity score, although reminiscent of the previously used margin score (Johansson et al.,, 2017; Linusson et al.,, 2018), uses the margin (distance to the maximum) on model output, not on the probability estimates. However, the interpretation as temperature calibration no longer applies, since softmax is not sparse. We also experiment with the non-conformity score (3.2) in §4.
@@ 4.1 Experimental Setup
We evaluate all approaches on tasks of varying difficulty: image classification on the CIFAR10, CIFAR100, and ImageNet datasets (Krizhevsky,, 2009; Deng et al.,, 2009) and text classification on the 20 Newsgroups dataset (Mitchell,, 1997). For each dataset, the original test data is split into calibration ($40\%$) and test sets ($60\%$) and the results reported are averaged over 5 random splits.
@@
For the CIFAR100 and ImageNet datasets, we finetune the vision transformer (ViT) model (Kolesnikov et al.,, 2021), obtaining average accuracies of approximately $0.86$ and $0.81$ on the test set, respectively. As for the 20 Newsgroups dataset, we finetune a BERT base model (Devlin et al.,, 2019) for sequence classification, obtaining an average test accuracy of $0.74$. Given the simplicity of the task on the CIFAR10 dataset, we train a convolutional neural network from scratch with final average test accuracy of $0.84$. More model evaluation details can be found in Appendix C.1.
@@
  • $\mathsf{RAPS}$: this is a modification of the standard procedure called regularized adaptive prediction sets (Angelopoulos et al.,, 2021), which makes use of two regularization parameters to improve the efficiency of the adaptive predictive sets (APS) predictors (Romano et al.,, 2020). In prior work, this has shown good efficiency and adaptiveness properties. Implementation of the $\mathsf{RAPS}$ procedure is done following the original paper [^7] Since using the same set to both calibrate the conformal predictor and find the optimal regularization hyperparameters would violate the exchangeability assumption, we split the original calibration data into two subsets for this procedure. (additional details can be found in Appendix B.1).
@@ 4.3 Adaptiveness
Given that conformal predictors are unable to provide conditional coverage theoretical guarantees (Vovk,, 2012), i.e., there is no lower bound on $\mathbb{P}\big{(}Y_{\text{test}}\in\mathcal{C}_{\alpha}(X_{\text{test}})|X_{\text{test}}\big{)}$ without extra assumptions, different measures have been proposed to calculate proximity to conditional coverage. Ideally, for any given partition of the data, we would have a coverage close to the $1-\alpha$ bound. Angelopoulos et al., (2021) introduced an adaptiveness criterion based on size-stratified coverage, by evaluating coverage in a partition based on the size of the predicted sets. Table 2 shows the size-stratified coverage for the ImageNet dataset, with $\alpha=0.01$ and $\alpha=0.1$. For $\alpha=0.1$, $\mathsf{log\text{-}margin}$ stands out from other methods, achieving a lower deviation from exact coverage on all set cardinalities. Size-stratified coverage improves for lower values of $\alpha$, and $\mathsf{InvProb}$, $\mathsf{log\text{-}margin}$ and $\mathsf{opt\text{-}entmax}$ exhibit the desired behavior—achieving expected coverage while still predicting small sets for easier examples.
@@ 5 DISCUSSION
The choice of non-conformity score plays a crucial role in the efficiency of a conformal predictor (Aleksandrova and Chertov,, 2021); consequently, finding useful model-agnostic scores is key for the adoption of conformal prediction as an actionable uncertainty quantification framework. In §3, we introduced a new family of non-conformity measures that we show to be competitive with standard scores on different classification tasks (§4). Expanding the range of possible non-conformity measures is particularly relevant, considering that the optimal score choice is known to be task-dependent (Laxhammar and Falkman,, 2010). Additionally, we show that these scores can be used to tune the adaptiveness of a conformal predictor.
@@ B.1 $\mathsf{RAPS}$
The $\mathsf{RAPS}$ procedure was implemented according to the original protocol introduced by Angelopoulos et al., (2021) by adapting the code provided by those authors. The method uses two hyperparameters to regularize the original adaptive prediction sets procedure (Romano et al.,, 2020): $\lambda_{\text{reg}}$ — a regularization penalty added to discourage inclusion in the prediction set, and $k_{\text{reg}}$ — the order from which classes receive the penalty term. We split the original calibration data into two sets: calibration data ($60\%$) and hyperparameter tuning data ($40\%$). For each split of the dataset, we find the pair of hyperparameters that minimizes the average prediction set size on hyperparameter tuning set, with $\lambda_{\text{reg}}\in\{0.001,0.01,0.1,1\}$ and $k_{\text{reg}}\in\{1,5,10,50\}$.
@@ C.4 Adaptiveness
In Figure 9 we can see the size-stratified coverage violation (SSCV) — introduced by Angelopoulos et al., (2021), it measures the maximum deviation from the desired coverage $1-\alpha$. Partitioning the possible size cardinalities into $G$ bins, $B_{1},...,B_{G}$, let $\mathcal{I}_{g}$ be the set of observations falling in bin $g$, with $g=1,...,G$, the SSCV of a predictor $C_{\alpha}$, for that bin partition is given by:

### v2

@@ 1 INTRODUCTION
The use of learned predictive models in many high-stakes applications (e.g., medical, legal, or financial) has stimulated extensive research on uncertainty quantification as a way to enhance predictions with reliable confidence estimates (Silva-Filho et al., 2023; Gawlikowski et al., 2023). In classification tasks, this corresponds to providing an estimate of the posterior probabilities of each of the classes, given the sample to be classified. In regression, this would correspond to providing, not only a point estimate, but also a confidence interval around that estimate.
@@
Unfortunately, modern deep networks are not well calibrated: the class probability estimates may be significantly different from the posterior probability, and thus cannot be trusted as confidence levels (Guo et al., 2017; Wenger et al., 2020). This observation has motivated considerable work on developing methods to obtain well-calibrated class probability estimates.
@@
To address this problem, conformal prediction2.1) has emerged as a powerful framework for uncertainty quantification that comes with strong theoretical guarantees under minimal assumptions, as it is model-agnostic and distribution-free (Vovk et al., 2005; Angelopoulos and Bates, 2023). Conformal predictors are set predictors, i.e., their outputs are either sets of labels (for classification tasks) or intervals (for regression tasks). Most common variants, namely split conformal prediction (Papadopoulos et al., 2002), do not require model retraining: prediction rules are created using a separate calibration set through the design of a non-conformity score, which is a measure of how unlikely an input-output pair is.
@@
A parallel research direction concerns sparse alternatives to the softmax activation (§2.2), i.e., capable of producing probability outputs where some labels receive zero probability, yet are differentiable, thus usable in backpropagation (Martins and Astudillo, 2016; Peters et al., 2019). The class of Fenchel-Young losses (Blondel et al., 2020) provides a framework to learn such sparse predictors, generalizing the softmax activation and the cross-entropy loss. An added benefit of sparse probability outputs is that the labels with nonzero probability can be directly interpreted as a set prediction, eliminating the need for creating these sets by thresholding probability values. Such set predictions have been used by Martins and Astudillo (2016) for multi-label classification, but not in the scope of uncertainty quantification.
@@
This paper connects the two lines of work above by conformalizing sparse activations (Figure 1; §3). We focus on the $\gamma$-entmax family [^1] Called $\alpha$-entmax by Peters et al. (2019). We choose $\gamma$ in this paper to avoid clash with the confidence level $\alpha$, commonly used in the conformal prediction literature. (Peters et al., 2019), which includes softmax ($\gamma=1$) and sparsemax ($\gamma=2$) as particular cases, and is sparse for $\gamma>1$. We create set predictions by including all labels receiving nonzero probability under those activations—these sets can be made larger or smaller by controlling a temperature parameter. We leverage the calibration techniques of conformal prediction to choose the temperature that leads to the desired coverage level.
@@ 2.1 Conformal Prediction
We focus on split conformal prediction, which does not involve model retraining (Papadopoulos et al., 2002). Given the predictor, $f$, and a collection of calibration samples, $\mathcal{D}_{\text{cal}}=\bigl{(}(x_{1},y_{1}),...,(x_{n},y_{n})\bigr{)}$, a conformal predictor is defined by a non-conformity score $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. The non-conformity score measures how unlikely an input-output pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is, compared to the remaining calibration data. Ideally, predictions $\hat{y}\in\mathcal{Y}$ yielding pairs $(x_{\text{test}},\hat{y})$ that are likely to occur in the data should have a low non-conformity score, and should thus be included in the prediction set $\mathcal{C}_{\alpha}(x_{\mathrm{test}})$.
@@
as proved by Vovk et al. (2005), without any other assumptions on the predictive model or the data distribution. Ideally, we would like the predictor with the required coverage to be efficient, i.e., to have a small average prediction set size $\mathbb{E}[|C_{\alpha}(X)|]$, and adaptive, i.e., instances that are harder to predict should yield larger prediction sets, representing higher uncertainty.
@@ 2.2 Sparse Activations
Since the exponential function is stricly positive, the softmax transformation never outputs any zero. Martins and Astudillo (2016) introduced an alternative transformation, sparsemax, which can produce sparse outputs while being almost-everywhere differentiable:
@@
As shown by Peters et al. (2019) and Blondel et al. (2020), sparsemax and softmax are particular cases of a family of $\gamma$-entmax transformations, defined as
@@
where $H_{\gamma}$ is a generalized entropy function (Tsallis, 1988), defined, for $\gamma>0$, as
@@
Peters et al. (2019) show that the solution of (5) has the form
@@
where $\mathbf{1}$ denotes a vector of ones, $\left[x\right]_{+}=\max\{x,0\}$, and $\tau$ is the (unique) constant ensuring normalization. A bisection algorithm for computing $\gamma$-entmax for general $\gamma$, as well as a more efficient algorithm for $\gamma=1.5$, have been proposed by Peters et al. (2019). We make use of these algorithms in our experiments in §4.
@@
Temperature scaling (Guo et al., 2017; Platt et al., 1999) is a simple and popular technique to calibrate the output probabilities from a predictive model. By multiplying the label scores $\bm{z}$ by a scaling factor $\beta>0$, which can be interpreted as an inverse temperature parameter, the probabilities resulting from the subsequent softmax transformation can be made more or less peaked. When $\beta=0$, $\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ becomes a uniform distribution, whereas in the limit $\beta\rightarrow+\infty$ (the “zero temperature limit”), $\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ tends to a one-hot vector, assigning probability 1 to the label with the largest score. [^4] If there are multiple labels tied with the same largest score, $\lim_{\beta\rightarrow+\infty}\operatorname*{\mathsf{softmax}}(\beta\bm{z})$ becomes a uniform distribution over those labels, assigning zero probability to all other labels.
Analogous limit cases apply also to $\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$, but there is a key qualitative difference (Martins and Astudillo, 2016; Blondel et al., 2020). For any $\gamma>1$, there is a finite threshold $\beta^{*}$ (corresponding to a “non-zero temperature”) above which $\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$ saturates as the zero temperature limit: $\beta\geq\beta^{*}\,\Rightarrow\,\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})=\lim_{\beta\rightarrow+\infty}\gamma\text{-}\operatorname*{\mathsf{entmax}}(\beta\bm{z})$. Increasing $\beta$, which may be seen as a regularization coefficient, from 00 to $\beta^{*}$ draws a regularization path where the support of $\gamma$-$\operatorname*{\mathsf{entmax}}(\beta\bm{z})$ varies from the full set of labels to a single label (assuming no ties), as shown in Figure 3. Therefore, using temperature scaling with $\gamma$-entmax, for $\gamma>1$, allows the sparsity of the output to be controlled. We exploit this fact in our proposed method described in §3.
@@ 3.2 Conformalizing $\gamma$-entmax
which is the log-odds ratio between the most probable class and the true one. Since the $\log$ function is monotonic, calibration of this non-conformity score leads to thresholding the odds ratio $p_{1}/p_{k(y)}$, which has an intuitive interpretation: labels whose probability is above a fraction of that of the most probable label are included in the prediction set. [^6] This non-conformity score, although reminiscent of the previously used margin score (Johansson et al., 2017; Linusson et al., 2018), uses the margin (distance to the maximum) on model output, not on the probability estimates. However, the interpretation as temperature calibration no longer applies, since softmax is not sparse. We also experiment with the non-conformity score (3.2) in §4.
@@ 4.1 Experimental Setup
We evaluate all approaches on tasks of varying difficulty: image classification on the CIFAR10, CIFAR100, and ImageNet datasets (Krizhevsky, 2009; Deng et al., 2009) and text classification on the 20 Newsgroups dataset (Mitchell, 1997). For each dataset, the original test data is split into calibration ($40\%$) and test sets ($60\%$) and the results reported are averaged over 5 random splits.
@@
For the CIFAR100 and ImageNet datasets, we finetune the vision transformer (ViT) model (Kolesnikov et al., 2021), obtaining average accuracies of approximately $0.86$ and $0.81$ on the test set, respectively. As for the 20 Newsgroups dataset, we finetune a BERT base model (Devlin et al., 2019) for sequence classification, obtaining an average test accuracy of $0.74$. Given the simplicity of the task on the CIFAR10 dataset, we train a convolutional neural network from scratch with final average test accuracy of $0.84$. More model evaluation details can be found in Appendix C.1.
@@
  • $\mathsf{RAPS}$: this is a modification of the standard procedure called regularized adaptive prediction sets (Angelopoulos et al., 2021), which makes use of two regularization parameters to improve the efficiency of the adaptive predictive sets (APS) predictors (Romano et al., 2020). In prior work, this has shown good efficiency and adaptiveness properties. Implementation of the $\mathsf{RAPS}$ procedure is done following the original paper [^7] Since using the same set to both calibrate the conformal predictor and find the optimal regularization hyperparameters would violate the exchangeability assumption, we split the original calibration data into two subsets for this procedure. (additional details can be found in Appendix B.1).
@@ 4.3 Adaptiveness
Given that conformal predictors are unable to provide conditional coverage theoretical guarantees (Vovk, 2012), i.e., there is no lower bound on $\mathbb{P}\big{(}Y_{\text{test}}\in\mathcal{C}_{\alpha}(X_{\text{test}})|X_{\text{test}}\big{)}$ without extra assumptions, different measures have been proposed to calculate proximity to conditional coverage. Ideally, for any given partition of the data, we would have a coverage close to the $1-\alpha$ bound. Angelopoulos et al. (2021) introduced an adaptiveness criterion based on size-stratified coverage, by evaluating coverage in a partition based on the size of the predicted sets. Table 2 shows the size-stratified coverage for the ImageNet dataset, with $\alpha=0.01$ and $\alpha=0.1$. For $\alpha=0.1$, $\mathsf{log\text{-}margin}$ stands out from other methods, achieving a lower deviation from exact coverage on all set cardinalities. Size-stratified coverage improves for lower values of $\alpha$, and $\mathsf{InvProb}$, $\mathsf{log\text{-}margin}$ and $\mathsf{opt\text{-}entmax}$ exhibit the desired behavior—achieving expected coverage while still predicting small sets for easier examples.
@@ 5 DISCUSSION
The choice of non-conformity score plays a crucial role in the efficiency of a conformal predictor (Aleksandrova and Chertov, 2021); consequently, finding useful model-agnostic scores is key for the adoption of conformal prediction as an actionable uncertainty quantification framework. In §3, we introduced a new family of non-conformity measures that we show to be competitive with standard scores on different classification tasks (§4). Expanding the range of possible non-conformity measures is particularly relevant, considering that the optimal score choice is known to be task-dependent (Laxhammar and Falkman, 2010). Additionally, we show that these scores can be used to tune the adaptiveness of a conformal predictor.
@@ B.1 $\mathsf{RAPS}$
The $\mathsf{RAPS}$ procedure was implemented according to the original protocol introduced by Angelopoulos et al. (2021) by adapting the code provided by those authors. The method uses two hyperparameters to regularize the original adaptive prediction sets procedure (Romano et al., 2020): $\lambda_{\text{reg}}$ — a regularization penalty added to discourage inclusion in the prediction set, and $k_{\text{reg}}$ — the order from which classes receive the penalty term. We split the original calibration data into two sets: calibration data ($60\%$) and hyperparameter tuning data ($40\%$). For each split of the dataset, we find the pair of hyperparameters that minimizes the average prediction set size on hyperparameter tuning set, with $\lambda_{\text{reg}}\in\{0.001,0.01,0.1,1\}$ and $k_{\text{reg}}\in\{1,5,10,50\}$.
@@ C.4 Adaptiveness
In Figure 9 we can see the size-stratified coverage violation (SSCV) — introduced by Angelopoulos et al. (2021), it measures the maximum deviation from the desired coverage $1-\alpha$. Partitioning the possible size cardinalities into $G$ bins, $B_{1},...,B_{G}$, let $\mathcal{I}_{g}$ be the set of observations falling in bin $g$, with $g=1,...,G$, the SSCV of a predictor $C_{\alpha}$, for that bin partition is given by: